OpenStudy (anonymous):

Please explain lim X ->2 (1/(x-2))- (1/|x-2|) = -infinity. I have worked out part of limit?

OpenStudy (abb0t):

Can you re-write it using LaTex? That's what the option is there for :) it makes it better to understand and avoid confusion when helping to answer a question.

OpenStudy (anonymous):

\[\lim_{x \rightarrow 2} (\frac{ 1 }{ x-2} - \frac{ 1 }{ \left| x-2 \right| })\] broke it nto tow parts to evalualte at 2 \[\lim_{x \rightarrow 2-} (\frac{ 1 }{ x-2} - \frac{ 1 }{ x-2})\] i boke i tup ad got ininfity - infinty \[\lim_{x \rightarrow 2+} (\frac{ 1 }{ x-2} - \frac{ 1 }{ -x+2 })\] ininfity + inifinty

OpenStudy (anonymous):

i used limit laws loosely, but this seems inaccurate

OpenStudy (abb0t):

For absolute, you need to evaluate it from both sides, when x > 2 and x <2

OpenStudy (abb0t):

and, as you know, when two limits do not approach the same from both sides, then the limit D.N.E

OpenStudy (anonymous):

i know the limit exists, because my answer is wrong. They are both negative infinity . I just am note sure where i went wrong

OpenStudy (amistre64):

when x>2, we get zero when x <2 go functional

OpenStudy (amistre64):

the limit from right is zero, the limit from the left is not 0

OpenStudy (amistre64):

f(x) = 2/(x-2) x < 0 f(x) = 0 x > 2

OpenStudy (amistre64):

lol, missed my piece domian ... x < 2 on top

OpenStudy (anonymous):

I don'tunderstand any of your post amistre64

OpenStudy (abb0t):

smh

OpenStudy (amistre64):

i dont know how to do the fancy schmancy piecewise in latex coeds ...

OpenStudy (amistre64):

it worst, create a table ....

OpenStudy (abb0t):

\[\begin{cases} & noob \text{ if } x>2 \\ & noob \text{ if } x<2 \end{cases}\]

OpenStudy (amistre64):

values of x bigger than 2; and values of x less than 2 you will see that a 2 sided limit does not exists. if your answer is wrong then look at the question and see if its asking for a 1 sided limit instead

OpenStudy (amistre64):

lol, now your just showing off

OpenStudy (abb0t):

http://www.codecogs.com/latex/eqneditor.php

OpenStudy (abb0t):

@amistre64

OpenStudy (anonymous):

Okay, apparently I type it in wrong when i went to check my answer or something. It does say now that the limit doesn't exist. Thanks for the help, I can only medal one, so please medal the other since you were both helpful to me.

OpenStudy (anonymous):

\[L.H.L=\lim_{x \rightarrow 2-}\left( \frac{ 1 }{ x-2 }-\frac{ 1 }{ \left| x-2 \right| } \right)\] put x=2-h,h>0,\[h \rightarrow0 asx \rightarrow2-\] \[L.H.L.=\lim_{h \rightarrow 0}\left( \frac{ 1 }{ 2-h-2 }-\frac{ 1 }{ \left| 2-h-2 \right| } \right)\] \[=\lim_{h \rightarrow 0}(\frac{ -1 }{h }-\frac{ 1 }{\left| -h \right| })=\lim_{H \rightarrow0}\left( -\frac{ 1 }{h }-\frac{ 1 }{ h }\right)\] \[=\lim_{h \rightarrow 0}\frac{ -2 }{h }\rightarrow-\infty \] \[R.H.L.=\lim_{x \rightarrow 2+}\left( \frac{ 1 }{x-2 }-\frac{ 1 }{ \left| x-2 \right| } \right)\] put x=2+h,h>0 ,\[h \rightarrow0as x \rightarrow2+\] \[R.H.L.=\lim_{h \rightarrow 0}\left( \frac{ 1 }{2+h-2 }-\frac{ 1 }{\left| 2+h-2 \right|} \right)\] \[=\lim_{h \rightarrow 0}\left( \frac{ 1 }{ h }-\frac{ 1 }{\left| h \right| } \right)=\lim_{h \rightarrow0}\left( \frac{ 1 }{h }-\frac{ 1 }{h } \right)=\lim_{h \rightarrow 0}=0\] \[L.H.L \neq R.H.L\] Limit does not exist.

OpenStudy (anonymous):

Nice.

OpenStudy (anonymous):

miss print 0 in the end.